Harvey Mudd College Mathematics Computing Laboratory

EVE EDT Tutorial

Written by:

R. Christopher Yoder

February 11, 1988

User Reference Document #1-88 Price: $0.25

1 Introduction

EVE EDT is an interactive text editor built using the VAX Tex Processing Utility (TPU). EVE EDT was based on the EVE (V 1.0-000) source, and it includes a keypad that provides nearly the same placement of keys and functionality of the EDT keypad. It also provides all of the functionality of EVE, along with multi-window capability from VTEDIT.TPU (by Ned Freed and Kevin Carosso), rectangular cut and paste routines that make a significant attempt to handle tabs, and several other nifty features too numerous to mention here. The keypad is as close to the EDT keypad as possible so that people who have “EDT fingers” can start editing right away. However, there are some slight differences which are listed in section titled “Changes from EDT” below.

See table 1 for the conventions followed in this document.

Notation  Meaning
<filename> Any valid filename, including its extension.
PFn  The row of function keys above a VT100 keypad.
KPn  A numerical key on the keypad (as opposed to the

keyboard). For example KP1 is the 1 key on the keypad.

Fn  The top row of function keys on a VT220 keyboard.
CTRL l  The control key and the letter l pressed together.
$ COMMAND A DCL command.

Table 1: Document conventions

2 Features

EVE EDT does not attempt to emulate EDT exactly. It simply provides a keypad that is similar to the one that EDT users know. However, there are a number of reasons why one should switch from EDT to EVE EDT. Following is a brief description of some of the major reasons to switch to EVE EDT.

First and foremost, EVE EDT is a TPU based editor. TPU is Digital’s answer to some of the fundamental problems in EDT. It is powerful, fast, and designed for the current state of computing.

TPU allows the user to display text on any size and shape of screen desired. EDT will only allow the user to use a screen that is 24 lines long and 80 characters wide or 24 lines long and 132 characters wide. People who have access to VAXstations or to terminals that allow more than 24 lines of text on the screen at one time cannot get EDT to display more than 24 lines at once on these devices, so TPU is a natural alternative.

TPU reads the entire file into memory at the start of the editing session so that moving around in the file goes much more quickly than in EDT. By storing the file in memory, certain operations such as searching for particular string of text goes much more quickly than it would in EDT.

Another nifty features provided by TPU includes the ability to “attach” to another process from a process running TPU. Allowing the user to attach to another process allows the user to have an editor session running in a subprocess and be able to go back and forth between the editing process and the main process. This is useful for debugging programs or putting those last touches on a document or any other applications where the user needs to make a number of small changes to a file, check the result, make more changes, etc. because there is much less overhead in switching from one process to another than there is in starting up an editor.

Beyond these basic features that TPU provides, EVE EDT provides a host of other features. Some of the more important include the ability to split any editing window, buffer and window manipulation commands, DCL command line compatability, and rectangular (regional) cut and paste, as well as some built in commands like “spell” and “sort” that execute the DCL command spell or sort on the entire buffer or selected region of text.

One feature that may need some explaining to those used to having only one window displayed on the screen at once is the concept of having more than one file or portion of the same file displayed. Splitting a window allows a user to display several files or different pieces of the same file at once. This type of feature is used for things like copying ideas from one file to another or for allowing a user to switch between two parts of the same file quickly. Besides the split window command, there are several window management commands that allow the user to change the size of each window to an appropriate size.

Typically, a user might split the current window, shrink one down a bit, get a new file to be edited (this automatically creates a new buffer) or switch buffers if the file is already in an editing buffer, and then go back to editing the other window. Whenever any operation that causes a portion of text to be stored in a buffer is performed, the reverse operation will work in whatever window you happen to be in. Thus a user can cut information in one window and then switch windows and insert the text into the new window. This has many applications including having a window into a buffer with an outline, another window with the document in it, and a third window with an old document where a similar problem was solved.

See the table 2 for some of the more important key definitions that are already made by EVE EDT.

3 Using EVE EDT

In order to use EVE EDT on the HMC Math Cluster to edit the file <filename>, the following command may be typed:

          $ EDIT/TPU/SECTION=PUBLIB:EVE EDTSECINI <filename>

If it is desired to make this TPU section file the default editor for all purposes, the following lines in the user’s LOGIN.COM are useful:

          $ DEFINE TPUSECINI PUBLIB:EVE EDTSECINI
          $ TPU == "EDIT/TPU"
          $ KTPU == "@PUBLIB:Kept TPU"
          $ DEFINE MAIL$EDIT "Callable TPU"

After making these definitions, every time that TPU is invoked to edit a file, either with the command $ EDIT/TPU <filename> or $ TPU <filename>, the EVE EDT section file will be used.

If an editing subprocess is desired, executing PUBLIB:Kept TPU.COM from DCL will either create a new editing subprocess if one does not exist or attach to an existing editor subprocess. From EVE EDT the command ATTACH or GOLD A will cause the user connect back to his main process. With the above definitions the command $ KTPU will execute PUBLIB:Kept TPU.COM.

Key  Definition

DCL Command Line Compatability

CTRL A  Insert/Overstrike Toggle
CTRL B  Recall Previous Command
CTRL E  Move to End of Current Line
Backspace  Move to Start of Current Line
Linefeed  Erase Previous Word
CTRL U  Erase to Start of Line

VT220 Keyboard Definitions

F10  Exit
F12  Start of Line (VT220 Backspace)
F13  Erase Previous Word (VT220 Linefeed)
F19  Write Modified
F20  Write File

Window Management

GOLD =  Split Window
GOLD 1  Single Window
GOLD –  Shrink Window
GOLD +  Expand Window
GOLD ↑  Previous Window
GOLD ↓  Next Window
GOLD ←  Shift Left 8
GOLD →  Shift Right 8

Rectangular Cut and Paste

GOLD .  Mark Corner
GOLD <  Extract Rectangle
GOLD >  Insert Rectangle
GOLD (  Copy Rectangle
GOLD )  Overlay Rectangle

Miscelaneous Definitions

GOLD A  Attach
GOLD B  Buffer

Table 2: Key Definitions

Finally, the definition of MAIL$EDIT above causes MAIL to use TPU and thus EVE EDT instead of EDT. This allows a user to have one standard editor for most editing tasks that may need to be done.

4 Changes from EDT

Those who use all of the keys on the EDT keypad should note the following exceptions to the EDT keypad placement in EVE EDT:

1. The REPEAT command: To repeat a command some number of times in EDT, a user would press the GOLD key (PF1), then type in the integer corresponding to the number of times for the command to repeat followed by the command.

Since EVE has a rather nice TPU procedure for repeating commands, it was tied to the GOLD PF2 key (PF1 and then PF2). To use it, press the GOLD (PF1) key and then the PF2 key. The user will be prompted for the number of times to repeat the next command. Following this the user will be prompted for the command to repeat.

2. The SUBS (GOLD ENTER) command: Since this command is serviced nicely by the REPLACE function (GOLD KP9) it wasn’t needed. In its place on the keyboard is the LEARN command.

Of the rest of the commands on the keypad, two have changed in the way that they are used and one in the functionality that is provided. These are:

1. The SPEC INS key (GOLD KP3) in EDT has been mapped to the QUOTE command: To insert a special character in EDT the user would press PF1 then enter the ASCII value of the character that was desired, then the PF1 key followed by the 3 key on the keypad (KP3). In EVE EDT this procedure is simplified so that now the PF1 (GOLD) key is pressed followed by KP3 followed by the character. (This approach means less fumbling with the ASCII charts to figure out what CTRL G or <ESC> maps to!) One thing to remember is that CTRL A maps to ASCII value 1, and CTRL Z maps to ASCII value 26.

2. The SEARCH command checks in the opposite direction if it can’t find the string searched for and then prompts the user to go to that occurrence of the string.

3. The REPLACE command uses the EVE REPLACE function: To replace one string with another in EDT the user would search for the string to replace, put the string to replace it with into the PASTE buffer, and then press PF1 followed by KP9. In EVE EDT the user starts off by pressing PF1 key followed by KP9. The REPLACE procedure then prompts for the old string followed the string that you want to replace it with. At each occurrence of the string the user will be prompted for whether or not to replace the string (yes/no), replace all of the strings found in the current direction (all), or quit replacing. REPLACE will also check in the opposite direction for any occurrences of the old string and prompt the user to go to that occurrence of the string.

Other than these differences, people who are used to using EDT should have very few problems adjusting to EVE EDT.